OpenStudy (anonymous):

Ashley has 100 books that she wants to give away at the rate of n books per week. Write a recursive function that represents the number of books Ashley has at any time. The recursive function that gives the number of books Ashley has at any time is

OpenStudy (tkhunny):

So, Week 0 gives 100 Week 1 gives 100 - n = Week 0 - n Week 2 gives 100 - 2n = Week 0 - 2n = Week 1 - n I'm seeing a pattern. You?